Solve the following differential equation subject to the given initial condition. (a)dy/d(theta)=ysin(theta); y(pi)=3. (b) x^2 dy/dx=y-xy; y(1)=1
1
Expert's answer
2022-02-23T15:08:33-0500

(a) "\dfrac{dy}{d\theta}=y\sin\theta; y(\pi)=3"


Solution:


Separate the variables

"\dfrac{dy}{y}=\sin\theta d\theta"

"\displaystyle\int\dfrac{dy}{y}=\int\sin\theta d\theta+c"

"\displaystyle\ln|y|=-\cos\theta+c"


General solution is

"\displaystyle y=e^{-\cos\theta+c}"


Rewrite general solution as

"\displaystyle y=e^{-\cos\theta}\cdot e^{c}"

To find "e^c" take initial conditions "y(\pi)=3" and substitute "\theta=\pi; y=3"

"\displaystyle 3=e^{-\cos\pi}\cdot e^{c}"

"\displaystyle 3=e^{-(-1)}\cdot e^{c}"

"\displaystyle 3=e\cdot e^{c}"

"\displaystyle e^{c}=\dfrac{3}{e}"

"\displaystyle y=e^{-\cos\theta}\cdot\frac{3}{e}"


Particular solution is

"\displaystyle y=3e^{-\cos\theta-1}"



(b) "x^2 \dfrac{dy}{dx}=y-xy; y(1)=1"


Solution:


Separate the variables

"x^2 \dfrac{dy}{dx}=y(1-x)"

"\dfrac{dy}{y}=\dfrac{(1-x)}{x^2}dx"

"\displaystyle\int\frac{dy}{y}=\int\frac{(1-x)}{x^2}dx+c"

"\displaystyle\int\frac{dy}{y}=\int\left(\frac{1}{x^2}-\frac{x}{x^2}\right)dx+c"

"\displaystyle\int\frac{dy}{y}=\int\left(\frac{1}{x^2}-\frac{1}{x}\right)dx+c"

"\displaystyle\ln|y|=-\frac{1}{x}-\ln|x|+c"

"\displaystyle\ln|y|+\ln|x|=-\frac{1}{x}+c"

"\displaystyle\ln|yx|=-\frac{1}{x}+c"

"\displaystyle yx=e^{-\frac{1}{x}+c}"


General solution is

"\displaystyle y=\frac{e^{-\frac{1}{x}+c}}{x}"


Substitute "x=1, y=1"

"\displaystyle 1=\frac{e^{-\frac{1}{1}+c}}{1}"

"\displaystyle 1=e^{-1+c}"

"\displaystyle e^0=e^{-1+c}"

"\displaystyle 0=-1+c"

"c=1"


Particular solution is

"\displaystyle y=\frac{e^{-\frac{1}{x}+1}}{x}"